BETTER MOUSETRAPS HAS DEVELOPED A NEW TRAP. IT CAN GO INTO PRODUCTION FOR AN INITIAL INVESTMENT IN EQUIPMENT OF 6 MILLION. THE EQUIPMENT WILL BE DEPRECIATED STRAIGHT-LINE OVER 6 YEARS TO A VALUE OF ZERO, BUT, IN FACT, IT CAN BE SOLD AFTER 6 YEARS FOR $500,000. THE FIRM BELIEVES THAT WORKING CAPITAL AT EACH DATE MUST BE MAINTAINED AT A LEVEL OF 10% OF NEXT YEAR'S FORECAST SALES. THE FIRM ESTIMATES PRODUCTION COST EQUAL TO $1.50 PER TRAP AND BELIEVES THAT THE TRAPS CAN BE SOLD FOR $4 EACH. SALES FORECASTS ARE YEAR 1 0.5, YEAR 2 0.6, YEAR 3 1.0, YEAR 4 1.0, YEAR 5 0.6, YEAR 6 0.2 AND THEREAFTER 0. THE PROJECT WILL COME TO AN END IN 5 YEARS WHEN THE TRAP BECOMES TECHNOLOGICALLY OBSOLETE. THE FIRM'S TAX BRACKET IS 35%, AND THE REQUIRED RATE OF RETURN ON THE PROJECT IS 12%. WHAT IS THE PROJECT NPV? WHAT IS THE IRR? WHAT IS PAYBACK IN YEARS?
Answer #1
1] [$ in millions] 0 1 2 3 4 5 6
Sales in millions of units 0.5000 0.6000 1.0000 1.0000 0.6000 0.2000
Sales revenue $          2.0000 $          2.4000 $           4.0000 $        4.0000 $        2.4000 $        0.8000
-Production cost $          0.7500 $          0.9000 $           1.5000 $        1.5000 $        0.9000 $        0.3000
-Depreciation $          1.0000 $          1.0000 $           1.0000 $        1.0000 $        1.0000 $        1.0000
=NOI $          0.2500 $          0.5000 $           1.5000 $        1.5000 $        0.5000 $      -0.5000
-Tax at 35% $          0.0875 $          0.1750 $           0.5250 $        0.5250 $        0.1750 $      -0.1750
=NOPAT $          0.1625 $          0.3250 $           0.9750 $        0.9750 $        0.3250 $      -0.3250
+Depreciation $          1.0000 $          1.0000 $           1.0000 $        1.0000 $        1.0000 $        1.0000
=OCF $          1.1625 $          1.3250 $           1.9750 $        1.9750 $        1.3250 $        0.6750
-Capital expenditure $          6.0000 Total
-Change in NWC $          0.2000 $          0.0400 $          0.1600 $                    -   $      -0.1600 $      -0.1600 $      -0.0800 $                 -  
+After tax salvage value = 0.5*(1-35%) $        0.3250
=FCF $        -6.2000 $          1.1225 $          1.1650 $           1.9750 $        2.1350 $        1.4850 $        1.0800 $        2.7625
PVIF at 12% [PVIF = 1/1.12^t] 1 0.89286 0.79719 0.71178 0.63552 0.56743 0.50663
PV at 12% $        -6.2000 $          1.0022 $          0.9287 $           1.4058 $        1.3568 $        0.8426 $        0.5472 $      -0.1166
NPV $        -0.1166
2] IRR is that discount rate for which NPV is 0. It has to be found out by trial and error.
Discounting at 11%:
PVIF at 11% 1 0.90090 0.81162 0.73119 0.65873 0.59345 0.53464
PV at 11% $        -6.2000 $          1.0113 $          0.9455 $           1.4441 $        1.4064 $        0.8813 $        0.5774 $        0.0660
IRR lies between 11% and 12%.
By simple interpolation IRR = 11%+1%*0.0660/(0.0660+0.1166) = 11.36%
3] Cumulative cash flows $        -6.2000 $        -5.0775 $        -3.9125 $         -1.9375 $        0.1975 $        1.6825 $        2.7625
Payback period = 3+1.9375/2.1350 = 3.91 Years
4] Recommendation:
As the NPV is negative, the project should not be undertaken.
Note:
The details regarding sales are given for 6 years. But, in the last portion it is said the trap will become obsoleted in 5 years.
The period of the project is taken as 6 years in line with the sales figures. The salvage value of the machine is also at EOY 6.